# Understanding LLM Agent Behaviours via Game Theory: Strategy Recognition, Biases and Multi-Agent Dynamics

### v1

@@ 3.2.1 Cooperation Rates Across Multiplication Factors
Figure 4: Comparison of cooperation rate evolution across 10 rounds for three LLM architectures under different multiplication factors ($r\in\{1.1,2.0,2.9\}$) and language conditions.
@@
Across all conditions, cooperation increases monotonically with the multiplication factor. The ordering $r=2.9>r=2.0>r=1.1$ holds uniformly across all LLM architectures (Claude vs. Mistral vs. ChatGPT) and languages (English vs. Vietnamese) for the majority of early, mid, and late rounds.
@@ 3.2.2 Cooperation Dynamics and End-Game Effects
Figure 5 examines how the coopera dynamics of play evolve over the ten rounds and whether there is any special instability at the end of the game. Because mismatch is computed from whether all three agents select the same action (100% match) or whether only two align ($\approx$ 66% match), the curves quantify the degree to which agents act coherently versus independently. Both panels plot different notions of “mismatch”: panel (a) tracks how often agents’ strategies differ from one another (strategy mismatch rate), while panel (b) breaks this down further by personality condition (cooperative vs. selfish) and looks at average coordination mismatch.
The trajectories in panel (a) show a consistent pattern across all models. By the later rounds, they all demonstrate substantially lower mismatch, indicating that the agents converge toward a common behavioral mode. Mistral Large exhibits the strongest convergence, with mismatch approaching zero by round 10, whereas Claude 3.5 Haiku and GPT-4o maintain slightly higher but still reduced levels of disagreement. This means that although cooperation levels collapse in the final rounds (as shown earlier in Figure 4), this collapse is not accompanied by chaotic or divergent play; instead, the agents converge together toward broadly similar non-cooperative choices.
Panel (b) further decomposes these coordination dynamics by agent personality. Under selfish personality instructions, all three models rapidly settle into a mutually consistent pattern: mismatch drops sharply after the first round and approaches zero well at the end of the game, with very little variation across rounds. Under cooperative instructions, however, mismatch remains substantially higher and decreases only gradually. The cooperative condition thus sustains a wider diversity of behaviors: some agents continue contributing while others defect, even deep into the game. Among the three models, Claude 3.5 Haiku displays the most persistent heterogeneity under cooperative instructions, while Mistral Large shows the fastest convergence in both personality conditions, and GPT-4o again falls between the other two in terms of both the rate and stability of alignment.
3.2.3 Model-Specific Behavioral Biases
Our cross-lingual and personality manipulation experiments reveal systematic behavioral biases inherent to each LLM architecture (Figure 6). These biases persist despite explicit personality framing, suggesting that model-specific training alignment strongly influences strategic decision-making in multi-agent settings.
@@ 3.2.3 Model-Specific Behavioral Biases
Figure 6: Model-specific behavioral biases across personality conditions and linguistic contexts. (a-b) Cooperation rates under selfish and cooperative personality prompts for English and Vietnamese. (c) Multi-dimensional comparison of behavioral characteristics across models.
@@
Figure 6 reveals three distinct behavioral profiles that emerge consistently across experimental conditions. Claude 3.5 Haiku exhibits the strongest prosocial bias, maintaining baseline cooperation even under explicit selfish framing. In the selfish scenario (Figure 6(a)), Claude achieves approximately 2% cooperation for both English and Vietnamese conditions, while other models approach or reach zero. This residual cooperation under selfish instructions suggests that Claude’s alignment training embeds prosocial tendencies that partially resist countervailing prompts. In the cooperative scenario (Figure 6(b)), the gap between languages indicates moderate sensitivity to linguistic framing. The large error bars visible across conditions reflect high behavioral variance, consistent with the elevated internal variability metric shown in the radar chart (Figure 6(c)).
@@
GPT-4o demonstrates the strongest adherence to personality instructions combined with extreme linguistic sensitivity. In the selfish scenario, GPT-4o exhibits zero cooperation across both languages, strictly following the assigned selfish persona without the prosocial leakage observed in Claude and Mistral. However, in the cooperative scenario, GPT-4o displays the most pronounced cross-lingual divergence among all models, with English cooperation substantially exceeding Vietnamese. This pattern is further confirmed by the radar chart, where GPT-4o occupies the extreme position along the cross-language inconsistency axis. The combination of perfect personality adherence in the selfish condition and dramatic linguistic effects in the cooperative condition suggests that GPT-4o’s behavior is multiplicatively determined by explicit instructions and implicit cultural associations, with neither factor independently dominating the other.
@@
Mistral Large presents a profile characterized by linguistic stability and moderate personality adherence. In the selfish scenario, Mistral achieves near-zero cooperation (approximately 1%) for both English and Vietnamese, demonstrating strong but not absolute adherence to selfish instructions. In the cooperative scenario (Figure 6(b)), Mistral maintains virtually identical cooperation rates across languages, exhibiting minimal cross-lingual variance. The radar chart confirms this language-invariant behavior, with Mistral scoring lowest on cross-language inconsistency while maintaining moderate levels across other dimensions. The consistently small error bars indicate deterministic response patterns with low behavioral variance. This stability makes Mistral’s behavior highly predictable across linguistic contexts, though the model exhibits lower baseline cooperation compared to Claude even under cooperative framing.
@@
The radar chart in Figure 6(c) synthesizes these behavioral signatures, illustrating how each model occupies a distinct region in the multidimensional strategy space. These findings indicate that model selection constitutes a strategic choice in multi-agent system design, as each architecture presents inherent trade-offs between cooperation bias, linguistic sensitivity, behavioral variance, and instruction adherence. Deployment decisions should explicitly account for these systematic biases rather than assuming models function as neutral strategic actors capable of arbitrary behavior through prompting alone.
3.3 Understanding Behavior LLMs by classification model
@@
We applied our trained supervised learning models to classify the behavioral strategies of LLM agents across the simulated games generated by the FAIRGAME framework. The analysis yields critical insights into the capabilities of machine learning for intent recognition and the intrinsic behavioral nature of Large Language Models.
@@
We first evaluated the robustness of different classifier architectures against execution noise, which simulates the stochasticity and potential "hallucinations" of LLMs. As shown in Figure 8, while Logistic Regression (LR) and Random Forest (RF) models achieved near-perfect accuracy (greater than 0.9) on clean data, their performance degraded when introduced to $5\%$ execution noise. In contrast, the Long Short-Term Memory (LSTM) network maintained the highest accuracy ($\sim 94\%$). This superiority stems from the LSTM’s recurrent architecture, which allows it to learn the sequential "context" of a strategy, effectively "forgiving" random deviations to identify the core behavioral pattern. Traditional models, which treat features as flattened vectors, lose this temporal coherence and are thus less resilient to the noise inherent in LLM outputs.
@@
In this study, we employed a selective filtering approach to ensure the reliability of our LLM behavioral strategy analysis. Specifically, we focused our analysis on game instances where the predicted strategy labels for both agents exhibited prediction probabilities exceeding 0.9 (90% confidence threshold). The decision to use high-confidence predictions is grounded in several key considerations:
  • Pattern Alignment with Theoretical Strategies: Samples with prediction probabilities above 0.9 indicate that the observed behavioral sequences of LLMs closely align with the canonical patterns defined by the four classical strategies (ALLD - Always Defect, ALLC - Always Cooperate, WSLS - Win-Stay Lose-Shift, and TFT - Tit-for-Tat). This strong correspondence suggests that these LLM behaviors can be meaningfully interpreted through the lens of established game-theoretic frameworks.
@@
  • Statistical Reliability: By focusing on high-confidence predictions, we minimize the risk of misclassification and ensure that our strategy distribution analysis reflects genuine behavioral patterns rather than classification artifacts. This threshold effectively filters out ambiguous cases where LLM behavior may represent transitional states, mixed strategies, or inconsistent play patterns.
@@
While our LSTM model demonstrates strong performance in strategy classification, they were originally designed as single-label classifiers among 4 strategies rather than multi-label with binary classification (e.g: TFT: 1, ALLD: 1, ALLC: 0, WSLS: 0). This architecture choice, while computationally efficient, introduces a limitation: when a behavioral pattern exhibits characteristics consistent with multiple strategies simultaneously, the model is constrained to output only the single most probable label. This limitiation is particularly attributed to the insufficient observation, which make pattern between strategies become ambiguous.
Hence, to address this constraint and ensure comprehensive coverage of LLM behavioral analysis, we adopt a hybrid labeling approach that combines model predictions with rule-based strategy assignment. The rule-based are clarified in Appendix … Specifically, in addition to utilizing the model’s predicted labels, we apply deterministic rule-based algorithms to identify and assign all potential strategy labels that align with the observed behavioral sequence, giving us the completeness of analysis and pattern coverage.
@@
Figure 8: This figure presents the strategic behavioral distribution across four Large Language Models (LLMs) in iterated Prisoner’s Dilemma gameplay. The analysis is based on high-confidence predictions (with probability greater than 0.9) from our trained classification model, with composite strategies expanded into individual pure strategies. Each pie chart represents the strategy distribution for Agent 1 across all game instances for a specific LLM model, providing a comparative view of inherent strategic preferences embedded within different language models.
The strategic preference analysis reveals significant heterogeneity in decision-making paradigms across LLM architectures, suggesting that strategic behavior is not merely a function of model scale but rather emerges from fundamental differences in training methodologies and alignment procedures.
Claude 3.5 Sonnet exhibits a cooperative-dominant behavioral pattern, with ALLC (31.7%) and WSLS (29.6%) emerging as its two most frequent strategies, reflecting a strong inclination toward both cooperative and adaptive responses. Llama 3.1 405B Instruct, in contrast, is characterized by a pronounced preference for WSLS (46.5%), which represents the highest proportion of any single strategy across all evaluated models and indicates a clear emphasis on adaptive conditional behavior rather than pure cooperation or defection.
@@
Figure 9: Strategy distribution across languages for multiple large language models (Claude, Llama, Mistral, GPT-4o) and the aggregated overview. The figures reveal clear cross-linguistic differences: Arabic consistently exhibits the highest rate of ALLD and the lowest rate of ALLC, while English and Chinese show a strong preference for the WSLS strategy. French tends to be more cooperative (AC, TFT), and Vietnamese often ranks second in ALLD usage.
@@
A striking finding is the profound impact of the language of interaction on strategic choice, a phenomenon we term “Linguistic-Cultural Priming". Under identical game rules and payoff matrices, the linguistic medium acted as a latent variable governing the agents’ “rationality". Across all models and the aggregated distribution, we observe consistent cross-linguistic variations in strategic behavior: Arabic consistently shows the highest proportion of ALLD, followed by Vietnamese, indicating a strong tendency toward non-cooperative or competitive behavior in these languages. When considering the aggregate distribution, Arabic also exhibits the lowest ALLC rate, forming a sharp contrast with its high defect rates, whereas French and Chinese demonstrate relatively stronger cooperative tendencies. English and Chinese consistently favor WSLS across all models, suggesting that prompts in these languages elicit more adaptive, outcome-dependent strategic behavior. French additionally displays moderate to high TFT usage, reflecting a reciprocal and conditionally cooperative pattern, while Vietnamese and Chinese generally adopt TFT less frequently. French and Chinese also rank higher in unconditional cooperation (AC), compared with Arabic, which scores lowest in this category. Despite variations in absolute percentages, the relative ordering of languages remains remarkably stable across Claude, Llama, Mistral, GPT-4o, and the aggregate figure, indicating that language-driven effects are stronger than model-specific differences. Overall, a clear cultural-linguistic clustering emerges: Arabic and Vietnamese lean toward defect-heavy strategies; English and Chinese prefer adaptive WSLS; and French shows the most cooperative profile (AC + TFT). These patterns may reflect latent cultural priors embedded within the models’ training data.
4 Summary and Discussion
4.1 Summary

### v2

@@ 4.2.1 Cooperation Rates Across Multiplication Factors
Figure 4: Comparison of cooperation rate evolution across 10 rounds for three LLM models under different multiplication factors ($r\in\{1.1,2.0,2.9\}$) and language conditions.
@@
Across all conditions, the overall level of cooperation increases with the multiplication factor. This is consistent with game theoretical predictions in PGG [sigmund:2010bo, van2012emergence]. This observation holds for all LLM models (Claude vs. Mistral vs. ChatGPT) and languages (English vs. Vietnamese) for the majority of early, mid, and late rounds.
@@ 4.2.2 Cooperation Dynamics and End-Game Effects
Figure 5 examines how the cooperation evolves over the ten rounds and whether there is any special instability at the end of the game. Because mismatch is computed from whether all three agents select the same action (100% match) or whether only two align ($\approx$ 66% match), the curves quantify the degree to which agents act coherently versus independently. Both panels plot different notions of “mismatch”: panel (a) tracks how often agents’ strategies differ from one another (strategy mismatch rate), while panel (b) breaks this down further by personality condition (cooperative vs. selfish) and looks at average coordination mismatch.
The trajectories in panel (a) show a consistent pattern across all models. By the later rounds, they all demonstrate substantially lower mismatch, indicating that the agents converge toward a common behavioural mode. Mistral Large exhibits the strongest convergence, with mismatch approaching zero by round 10, whereas Claude 3.5 Haiku and GPT-4o maintain slightly higher but still reduced levels of disagreement. This means that although cooperation levels collapse in the final rounds (as shown earlier in Figure 4), this collapse is not accompanied by chaotic or divergent play; instead, the agents converge together toward broadly similar non-cooperative choices.
Panel (b) further decomposes these coordination dynamics by agent personality. Under selfish personality instructions, all three models rapidly settle into a mutually consistent pattern: mismatch drops sharply after the first round and approaches zero well at the end of the game, with very little variation across rounds. Under cooperative instructions, however, mismatch remains substantially higher and decreases only gradually. The cooperative condition thus sustains a wider diversity of behaviours: some agents continue contributing while others defect, even deep into the game. Among the three models, Claude 3.5 Haiku displays the most persistent heterogeneity under cooperative instructions, while Mistral Large shows the fastest convergence in both personality conditions, and GPT-4o again falls between the other two in terms of both the rate and stability of alignment.
4.2.3 Model-Specific behavioural Biases
Our cross-lingual and personality manipulation experiments reveal systematic behavioural biases inherent to each LLM architecture (Figure 6). These biases persist despite explicit personality framing, suggesting that model-specific training alignment strongly influences strategic decision-making in multi-agent settings.
@@ 4.2.3 Model-Specific behavioural Biases
Figure 6: Model-specific behavioural biases across personality conditions and linguistic contexts. (a-b) Cooperation rates under selfish and cooperative personality prompts for English and Vietnamese. (c) Multi-dimensional comparison of behavioural characteristics across models.
@@
Figure 6 reveals three distinct behavioural profiles that emerge consistently across experimental conditions. Claude 3.5 Haiku exhibits the strongest prosocial bias, maintaining baseline cooperation even under explicit selfish framing. In the selfish scenario (Figure 6(a)), Claude achieves approximately 2% cooperation for both English and Vietnamese conditions, while other models approach or reach zero. This residual cooperation under selfish instructions suggests that Claude’s alignment training embeds prosocial tendencies that partially resist countervailing prompts. In the cooperative scenario (Figure 6(b)), the gap between languages indicates moderate sensitivity to linguistic framing. The large error bars visible across conditions reflect high behavioural variance, consistent with the elevated internal variability metric shown in the radar chart (Figure 6(c)).
@@
GPT-4o demonstrates the strongest adherence to personality instructions combined with extreme linguistic sensitivity. In the selfish scenario, GPT-4o exhibits zero cooperation across both languages, strictly following the assigned selfish persona without the prosocial leakage observed in Claude and Mistral. However, in the cooperative scenario, GPT-4o displays the most pronounced cross-lingual divergence among all models, with English cooperation substantially exceeding Vietnamese. This pattern is further confirmed by the radar chart, where GPT-4o occupies the extreme position along the cross-language inconsistency axis. The combination of perfect personality adherence in the selfish condition and dramatic linguistic effects in the cooperative condition suggests that GPT-4o’s behaviour is multiplicatively determined by explicit instructions and implicit cultural associations, with neither factor independently dominating the other.
@@
Mistral Large presents a profile characterized by linguistic stability and moderate personality adherence. In the selfish scenario, Mistral achieves near-zero cooperation (approximately 1%) for both English and Vietnamese, demonstrating strong but not absolute adherence to selfish instructions. In the cooperative scenario (Figure 6(b)), Mistral maintains virtually identical cooperation rates across languages, exhibiting minimal cross-lingual variance. The radar chart confirms this language-invariant behaviour, with Mistral scoring lowest on cross-language inconsistency while maintaining moderate levels across other dimensions. The consistently small error bars indicate deterministic response patterns with low behavioural variance. This stability makes Mistral’s behaviour highly predictable across linguistic contexts, though the model exhibits lower baseline cooperation compared to Claude even under cooperative framing.
@@
The radar chart in Figure 6(c) synthesizes these behavioural signatures, illustrating how each model occupies a distinct region in the multidimensional strategy space. These findings indicate that model selection constitutes a strategic choice in multi-agent system design, as each architecture presents inherent trade-offs between cooperation bias, linguistic sensitivity, behavioural variance, and instruction adherence. Deployment decisions should explicitly account for these systematic biases rather than assuming models function as neutral strategic actors capable of arbitrary behaviour through prompting alone.
4.3 Understanding behaviour LLMs by classification model
@@
We applied our trained supervised learning models to classify the behavioural strategies of LLM agents across the simulated games generated by the FAIRGAME framework. The analysis yields critical insights into the capabilities of machine learning for intent recognition and the intrinsic behavioural nature of LLMs.
@@
We first evaluated the robustness of different classifier architectures against execution noise, which simulates the stochasticity and potential "hallucinations" of LLMs. As shown in Figure 8, while Logistic Regression (LR) and Random Forest (RF) models achieved near-perfect accuracy (greater than 0.9) on clean data, their performance degraded when introduced to $5\%$ execution noise. In contrast, the Long Short-Term Memory (LSTM) network maintained the highest accuracy ($\sim 94\%$). This superiority stems from the LSTM’s recurrent architecture, which allows it to learn the sequential "context" of a strategy, effectively "forgiving" random deviations to identify the core behavioural pattern. Traditional models, which treat features as flattened vectors, lose this temporal coherence and are thus less resilient to the noise inherent in LLM outputs.
@@
In this study, we employed a selective filtering approach to ensure the reliability of our LLM behavioural strategy analysis. Specifically, we focused our analysis on game instances where the predicted strategy labels for both agents exhibited prediction probabilities exceeding 0.9 (90% confidence threshold). The decision to use high-confidence predictions is grounded in several key considerations:
  • Pattern Alignment with Theoretical Strategies: Samples with prediction probabilities above 0.9 indicate that the observed behavioural sequences of LLMs closely align with the canonical patterns defined by the four classical strategies (ALLD - Always Defect, ALLC - Always Cooperate, WSLS - Win-Stay Lose-Shift, and TFT - Tit-for-Tat). This strong correspondence suggests that these LLM behaviours can be meaningfully interpreted through the lens of established game-theoretic frameworks.
@@
  • Statistical Reliability: By focusing on high-confidence predictions, we minimize the risk of misclassification and ensure that our strategy distribution analysis reflects genuine behavioural patterns rather than classification artifacts. This threshold effectively filters out ambiguous cases where LLM behaviour may represent transitional states, mixed strategies, or inconsistent play patterns.
@@
While our LSTM model demonstrates strong performance in strategy classification, they were originally designed as single-label classifiers among 4 strategies rather than multi-label with binary classification (e.g: TFT: 1, ALLD: 1, ALLC: 0, WSLS: 0). This architecture choice, while computationally efficient, introduces a limitation: when a behavioural pattern exhibits characteristics consistent with multiple strategies simultaneously, the model is constrained to output only the single most probable label. This limitiation is particularly attributed to the insufficient observation, which make pattern between strategies become ambiguous.
Hence, to address this constraint and ensure comprehensive coverage of LLM behavioural analysis, we adopt a hybrid labeling approach that combines model predictions with rule-based strategy assignment. The rule-based are clarified in Appendix … Specifically, in addition to utilizing the model’s predicted labels, we apply deterministic rule-based algorithms to identify and assign all potential strategy labels that align with the observed behavioural sequence, giving us the completeness of analysis and pattern coverage.
@@
Figure 8: This figure presents the strategic behavioural distribution across four LLMs in iterated Prisoner’s Dilemma gameplay. The analysis is based on high-confidence predictions (with a probability greater than 0.9) from our trained classification model, with composite strategies expanded into individual pure strategies. Each pie chart represents the strategy distribution for Agent 1 across all game instances for a specific LLM model, providing a comparative view of inherent strategic preferences embedded within different language models.
The strategic preference analysis reveals significant heterogeneity in decision-making paradigms across LLM architectures, suggesting that strategic behaviour is not merely a function of model scale but rather emerges from fundamental differences in training methodologies and alignment procedures.
Claude 3.5 Sonnet exhibits a cooperative-dominant behavioural pattern, with ALLC (31.7%) and WSLS (29.6%) emerging as its two most frequent strategies, reflecting a strong inclination toward both cooperative and adaptive responses. Llama 3.1 405B Instruct, in contrast, is characterized by a pronounced preference for WSLS (46.5%), which represents the highest proportion of any single strategy across all evaluated models and indicates a clear emphasis on adaptive conditional behaviour rather than pure cooperation or defection.
@@
Figure 9: Strategy distribution across languages for multiple LLMs (Claude, Llama, Mistral, GPT-4o) and the aggregated overview. The figures reveal clear cross-linguistic differences: Arabic consistently exhibits the highest rate of ALLD and the lowest rate of ALLC, while English and Chinese show a strong preference for the WSLS strategy. French tends to be more cooperative (AC, TFT), and Vietnamese often ranks second in ALLD usage.
@@
A striking finding is the profound impact of the language of interaction on strategic choice, a phenomenon we term “Linguistic-Cultural Priming". Under identical game rules and payoff matrices, the linguistic medium acted as a latent variable governing the agents’ “rationality". Across all models and the aggregated distribution, we observe consistent cross-linguistic variations in strategic behaviour: Arabic consistently shows the highest proportion of ALLD, followed by Vietnamese, indicating a strong tendency toward non-cooperative or competitive behaviour in these languages. When considering the aggregate distribution, Arabic also exhibits the lowest ALLC rate, forming a sharp contrast with its high defect rates, whereas French and Chinese demonstrate relatively stronger cooperative tendencies. English and Chinese consistently favor WSLS across all models, suggesting that prompts in these languages elicit more adaptive, outcome-dependent strategic behaviour. French additionally displays moderate to high TFT usage, reflecting a reciprocal and conditionally cooperative pattern, while Vietnamese and Chinese generally adopt TFT less frequently. French and Chinese also rank higher in unconditional cooperation (AC), compared with Arabic, which scores lowest in this category. Despite variations in absolute percentages, the relative ordering of languages remains remarkably stable across Claude, Llama, Mistral, GPT-4o, and the aggregate figure, indicating that language-driven effects are stronger than model-specific differences. Overall, a clear cultural-linguistic clustering emerges: Arabic and Vietnamese lean toward defect-heavy strategies; English and Chinese prefer adaptive WSLS; and French shows the most cooperative profile (AC + TFT). These patterns may reflect latent cultural priors embedded within the models’ training data.
5 Discussion
Our study presents several methodological constraints that limit the generalizability of findings. First, the ten-round game horizon across both Prisoner’s Dilemma and Public Goods Game experiments, while sufficient to observe initial strategic patterns and end-game dynamics, may be too short to capture sophisticated long-term behaviours such as reputation-building, conditional strategies, or forgiveness strategies that require extended interaction histories to stabilise. Prior work on human subjects shows that experiments under 25 rounds primarily capture the learning phase, not stabilised strategic behaviour, meaning that short sequences risk misrepresenting the true strategy dynamics [inferenceStrategies].
Second, our linguistic coverage is restricted to English and Vietnamese, preventing broader conclusions about how linguistic-cultural priming operates across diverse language families with different grammatical structures, collectivist-individualist orientations, or varying representation densities in LLM training corpora.
Third, our experimental design examines only two game-theoretic settings with limited parametric variation-the Prisoner’s Dilemma explores payoff magnitude scaling but maintains symmetric two-player structure, while the Public Goods Game uses a fixed group size of three agents, which represents the minimal configuration for non-dyadic effects but precludes examination of larger-group phenomena such as diffusion of responsibility or coalition formation. We will explore other settings of varying strategic natures such as coordination, trust and fairness [wang2015universal, Guth2014MoreLiterature].
Fourth, the machine learning-based intent recognition pipeline focuses exclusively on four canonical strategies (ALLC, ALLD, TFT, WSLS) derived from classical game theory, and our high-confidence filtering approach prioritizes interpretability while necessarily excluding behavioural patterns that may represent emergent hybrid strategies not captured by this taxonomy [inferenceStrategies].
Finally, the absence of parallel human behavioural experiments under matched conditions prevents rigorous assessment of whether the observed cross-lingual cooperation gaps, model-specific prosocial biases, and strategic patterns genuinely reflect human-like reasoning or constitute artifacts of model-specific alignment procedures, limiting our ability to evaluate the ecological validity of LLM agents in real-world multi-agent scenarios.
Future work will focus on advancing strategy identification to understand how LLM agents behave in repeated social dilemmas and whether these behaviours differ from those of humans, using game-playing trajectories as the basis for comparison. We will collect substantially more LLM data by extending the original FAIRGAME 10-round interactions into long repeated round games (e.g. with over 100 rounds). This would increase observations for downstream analysis and expand contextual information by incorporating both action histories and Chain-of-Thought (CoT) [NEURIPS2022_9d560961] traces into strategy inference. With such larger datasets, we will explore more efficient strategy-identification techniques, such as clustering on extended trajectories, clustering on CoT sequences combined with actions, supervised models beyond logistic regression and LSTMs, and unsupervised methods including Hidden Markov Models [inferenceStrategies], in order to support both classification and discovery of emergent strategies.
We also plan to benchmark these inferred strategies against human behavioural datasets to analyse similarities and divergences between LLM and human strategic behaviour, and pursue new directions such as multi-player games beyond two-player games. Finally, we aim to extend the FAIRGAME setup by experimenting with varied prompt templates, adding more languages beyond English and Vietnamese, enabling dynamic-importance or continual-learning configurations, mixing different games within multi-round sessions to test whether strategy shifts occur, and optionally allowing inter-agent communication to examine how environmental and linguistic factors shape LLM strategy formation.
6 Conclusion